### v1

# OrionNav: Online Planning for Robot Autonomy with Context-Aware LLM and Open-Vocabulary Semantic Scene Graphs

Abstract
Enabling robots to autonomously navigate unknown, complex, dynamic environments and perform diverse tasks remains a fundamental challenge in developing robust autonomous physical agents. They must effectively perceive their surroundings while leveraging world knowledge for decision-making. While recent approaches utilize vision-language and large language models for scene understanding and planning, they often rely on offline processing, external computing, or restrictive environmental assumptions. We present a novel framework for efficient and scalable real-time, onboard autonomous navigation that integrates multi-level abstraction in both perception and planning in unknown large-scale environments that change over time. Our system fuses data from multiple onboard sensors for localization and mapping and integrates it with open-vocabulary semantics to generate hierarchical scene graphs. An LLM-based planner leverages these graphs to generate high-level task execution strategies, which guide low-level controllers in safely accomplishing goals. Our framework’s real-time operation enables continuous updates to scene graphs and plans, allowing swift responses to environmental changes and on-the-fly error correction. This is a key advantage over static or rule-based planning systems. We demonstrate our system’s efficacy on a quadruped robot navigating large-scale, dynamic environments, showcasing its adaptability and robustness in diverse scenarios.
I INTRODUCTION
Autonomous physical agents operating in real-world environments require robust and scalable scene representations capable of adapting to continuously evolving conditions, enabling them to perform complex tasks such as navigation, manipulation, and interaction [1, 2, 3, 4]. Moreover, these agents must be capable of swiftly recognizing when a plan’s execution is unsuccessful and dynamically adjusting their strategies, ensuring resilience in the face of unexpected changes in the environment. This necessitates integrated approaches to both perception and planning, where agents must recognize and classify objects, construct structured models of their surroundings, and formulate multi-step plans to make informed decisions and execute tasks. A fundamental challenge lies in developing these representations and plans in real-time using onboard sensors while ensuring adaptability to the dynamically changing nature of real-world settings.
Simultaneous Localization and Mapping (SLAM) has long been the cornerstone for enabling robots to build maps and localize themselves within them [5, 6, 7]. While traditional SLAM techniques primarily focus on geometric and spatial information, they lack the semantic understanding crucial for higher-level reasoning and task execution. Object detection and segmentation pipelines have been previously integrated into these systems to bridge this gap, allowing them to identify and categorize objects within their environment while also refining the map and odometry [8, 9, 10]. Recent advancements in open-vocabulary detection and segmentation [11] are now being integrated into the mapping process, enhancing the robot’s environmental understanding by identifying and classifying objects within varying environments without requiring retraining [12]. However, these approaches often rely on pre-built semantic maps and struggle with real-time operation, limiting their applicability in dynamically changing, large-scale cluttered environments. Effective navigation and task performance in such settings demand scene representations organized hierarchically across multiple abstraction levels. Scene graphs have emerged as a powerful framework to meet this need, capturing objects and their interrelationships in a structured manner [13, 14, 15, 16]. These graphs not only encode geometric information but also capture semantic attributes and relations, providing a concise depiction of the scene to facilitate complex planning using large language models (LLMs) to achieve a desired goal [17, 18, 19, 20, 16].
Figure 1: Overview of our OrionNav framework. The OrionNav system fuses data from onboard LiDAR and odometry sensors for robust localization and mapping, while integrating open-world semantics to produce a semantic object map of the environment. This map is then clustered into distinct rooms, and room labels are assigned using the Llama3 LLM, generating a hierarchical scene graph. An LLM-based planner utilizes this scene graph, along with user commands, to create a high-level task execution plan, guiding low-level controllers to safely and efficiently achieve designated goals.
The advent of LLMs has transformed autonomous planning due to their ability to understand and reason about complex, context-rich scenarios [21, 22, 23, 24, 25, 17, 18, 19, 20, 26]. Thus, our system integrates LLMs with hierarchical scene graph representations, enabling real-time adjustments and error corrections based on the agent’s evolving understanding of its environment. This enables the generation of context-aware high-level plans for autonomous task execution using information from the scene graphs. The execution of these high-level plans is aided by lower-level planners and controllers. We demonstrate this on a quadruped robot, focusing on navigation tasks that involve reaching specific objects as directed by the user, in unknown environments that change over time. This requires robust coordination between high-level planning and lower-level motion control, including obstacle avoidance. This coordination is crucial for effectively translating complex plans into precise physical actions, allowing the robot to navigate and interact with its surroundings in real time.
The quadruped robot utilizes a combination of RGB-D and LiDAR sensors to autonomously navigate an unknown environment by first generating an object map through integrated open-vocabulary semantic segmentation and SLAM. The maps are continuously updated in real time as the robot navigates around the scene. The objects in the map are spatially clustered and associated with specific rooms using an onboard LLM to construct a hierarchical scene graph. This graph representation enables an LLM to formulate a contextually appropriate plan for executing user-defined navigation tasks. A key innovation of our work lies in the system’s ability to continuously update the scene representations and plan online during robot operation. This allows the robot to swiftly respond to environmental changes and correct planning errors on the fly. This capability is crucial for maintaining effectiveness in dynamic, unpredictable real-world settings where static pre-programmed behaviors often fail. The high-level plan guides the low-level motion planner to generate goal-directed navigation commands, which are then refined by an obstacle-avoiding control barrier function (CBF) controller. These refined commands ensure safe and efficient navigation toward the designated goals. The overall flow of the proposed framework is illustrated in Fig. 1. All modules except for the LLM used for planning operate onboard the robot in real-time while demonstrating autonomy in complex environments to perform diverse tasks. Our key contributions are:
  • A framework with multi-level abstractions of perception and planning for autonomous language-driven navigation that adapts to dynamic environments and task failures.
  • A robust and efficient open-vocabulary 3D semantic mapping approach to generate and update hierarchical scene graphs during robot operation.
  • Integration of 3D semantic scene graphs with LLMs for high-level planning that guides lower-level motion planners and CBF-based obstacle avoidance.
  • Implementation of our framework on a quadruped to perform various tasks autonomously in real-world large-scale environments utilizing onboard sensors and computing.
The rest of the paper is organized as follows: Section II covers related work, Section III presents the problem and Section IV describes our framework, Section V shows experimental results, and Section VI concludes the paper.
Figure 2: Our method makes use of the semantic constructs of indoor environments to generate a hierarchical scene graph from the semantic object map. The objects are then clustered based on the density of the objects in the scene. An LLM is then queried with the labels of the objects in each cluster and a set of candidate room labels. Finally, the generated hierarchical graph is passed to the LLM-based planner.
II RELATED WORKS
II-A Scene Graph
Scene Graphs have proven to be an efficient and compact way of representing any given environment. This is achieved by using a tree structure with defined nodes representing compressed semantic features derived from high-dimensional map data obtained through Semantic SLAM. Additionally, edges are constructed between node pairs to represent relative spatial or semantic constraints. Such graphs aid in understanding global and local characteristics of the map, which in turn helps with informed and accurate planning. The map data represented by the nodes can be image instances, object instances conveyed through 3D point clouds [27, 13, 28], 2D masks [24], or larger features like annotated spatial regions [29]. To connect these features with respect to each other, edges are coded to represent geometric or semantic relations between the nodes, such as apple <on> table [30]. Such structures have been commonly used for image retrieval [31, 32], object retrieval [16], and task and motion planning [33].
Earlier works have relied on Conditional Random Fields (CRFs) [34] to model graph nodes and edges. Advancements in deep learning have enabled the training of Graph Neural Networks (GNNs) to generate scene graphs from image instances and object masks [35, 36]. However, the advent of Large Visual-Language Models has facilitated the creation of open-vocabulary annotations of nodes and the appending of language-aligned features to improve the performance of downstream tasks. Foundational models [37, 38, 39, 40] are widely used to enrich scene graph nodes with semantic features such as CLIP [37] embeddings and linguistic annotations like labels, captions, and descriptions. Many methods [13, 14, 15, 16] associate object nodes with corresponding CLIP features obtained by fusing individual embeddings from multiple 2D views, masked views, or re-projected 3D features. Works like ConceptGraphs augment the nodes with language captions to enable the use of generated scene graphs with LLM-based planners. Along with node enhancement, foundational models are also capable of generating semantic relations between nodes. VLMs such as LLaVa [39] or BLIP [41] are used either to directly generate spatial semantic relations between objects using SLAM-derived poses and 2D camera views [24], or indirectly to train graph networks to populate the node edges [13]. However, most of these approaches are computationally intensive, preventing online scene graph generation. This limits the applicability of such works to static environments. Our method, on the other hand, realizes a novel scene graph representation and generation, enabling it to work online.
The works closest to our approach are HOV-SG [16] and Clio [15], which is an extension of their previous closed-set scene graph generation method called Hydra [42]. HOV-SG segments the scene using the Segment Anything Model [43], projects the segments as point clouds, fuses CLIP-segment features, and clusters them to generate 3D masks. These are used to generate a scene graph for indoor environments [44, 45], where nodes are characterized using fused CLIP features and heuristics to generate a scene graph with a hierarchy of buildings, floors, and rooms. However, their computationally expensive clustering and classification methods make it difficult to implement in real time. On the other hand, Clio attempts efficient graph generation using the Information Bottleneck framework. By clustering the scene into task-conditioned primitives, they limit the scene graph to task-relevant semantic concepts, leading to efficient scene graph generation in real time for object retrieval. However, such scene graphs are applicable only to a predefined task list, preventing usage on out-of-list tasks. Our method aims to mitigate the shortcomings of both these methods while enabling real-time scene graph generation and execution of various tasks.
II-B Semantic Mapping
Scene segmentation is a fundamental task in computer vision that aims to partition an image or a 3D scene into semantically meaningful regions, playing a crucial role in applications such as object recognition, scene understanding, and robot navigation [46, 47, 2, 4, 48, 49, 50]. Deep learning has revolutionized this field, enabling the development of highly accurate and efficient models across various paradigms. These include semantic segmentation (inferring high-level category concepts), instance segmentation (grouping foreground pixels into distinct object instances), panoptic segmentation (combining semantic and instance segmentation), and entity segmentation (inferring masks of object instances without classification) [51, 52, 11]. Panoptic segmentation, in particular, predicts both class and instance labels for each pixel, assigning instance labels to foreground objects (‘things’) and single labels to background parts (‘stuff’). The advent of foundation vision-language models like CLIP [37] has enabled more sophisticated approaches to fine-grained recognition and segmentation of objects in scenes in an open vocabulary setting [12]. These models leverage large-scale pre-training on diverse image-text pairs, allowing them to generalize to a wide range of visual concepts and enabling zero-shot learning capabilities which are made robust with out-of-distribution detection [53]. Recent developments have further extended these capabilities to open-vocabulary semantic and instance segmentation tasks, demonstrating impressive zero-shot generalization to novel object categories [11, 54, 55, 56, 57, 58, 59, 60].
Reconstruction of 3D structure with object recognition and grouping remains a fundamental challenge in computer vision, especially in the context of 3D scene understanding. These methods help a large language model help understand the scene through reasoning [61, 62] for building grounded 3D large language models [63, 64, 65, 66, 67, 68, 69]. The semantic information aids in both understanding the scene as well as improving the accuracy of the map and odometry for semantic SLAM [8, 70, 71, 72, 73, 74, 75, 76, 77, 78]. End-to-end 3D methods that predict 3D semantic information directly from sensor modalities like camera and depth, or from intermediate representations like reconstructed point clouds or voxel-based representations, can capture complex 3D relationships. Recent works have significantly improved the performance of open vocabulary 3D segmentation using point clouds from LiDAR or depth data from RGB-D cameras [79, 80, 81, 82, 83, 84, 85, 86, 87, 88]. In addition, there have also been approaches fusing embeddings from open world vision language models like CLIP with neural radian fields, Gaussian splats or implicit neural network based representation [29, 89, 90, 91] for generating a consistent segmentation of the scene. The open vocabulary allows for greater flexibility in real-world applications and can help an LLM to understand the scene with multiple objects, which multi-modal LLMs hallucinate or fail to perceive [92]. However, these methods do not scale well in larger scene making them computationally intensive and memory-consuming, rendering them impractical for real-time segmentation in robotic applications.
OrionNav draws inspiration from prior works while prioritizing efficiency, striking a balance between adaptability and real-time performance. The proposed method leverages the strengths of both 2D and 3D approaches, utilizing the efficiency and generalization capabilities of 2D CLIP-based models while incorporating 3D geometric information to enhance spatial understanding. It integrates a CLIP-based open vocabulary 2D segmentation model with 3D geometric information from map and odometry for improved scene understanding. We utilize the FC-CLIP [93] model with a ConvNeXt [94] CLIP [37] backbone, enabling efficient processing of high-resolution images where transformer-based models struggle. This architecture maintains translation equivariance and allows multi-scale feature fusion, crucial for detailed environmental segmentation in robotic applications. The system performs lightweight updates to the semantic object map leveraging spatial and semantic properties of the scene and observed objects in it. This map can be further utilized by a scene graph to generate hierarchical object clusters and relationships. This hybrid approach offers a robust, efficient solution that addresses trade-offs between computational efficiency and comprehensive 3D scene understanding.
II-C Multi-modal Language Models for Robotic Task Planning
Robotic task planning research, in recent years, has benefitted from the use of LLMs [21, 22, 23, 25, 17, 18, 19, 20, 26]. Early studies in this field primarily focused on manipulation tasks and humanoid simulations [21, 22, 23]. Recent works have shifted toward generating plans for mobile robots operating in larger environments [17, 18, 19, 20, 26, 95]. Strategies such as environment feedback through the use of the Python assert keyword in the plans [21], affordance functions [25], and success detection and scene description [23] have been used to ground the LLM planners in the environment. Studies focusing on mobile robots operating in large environments often utilize scene graphs to ground LLMs in their environment [17, 18, 19, 20, 24, 16], a strategy we also adopt in this framework.
ConceptGraphs [24] and HOV-SG [16] use LLM to retrieve the target object from the prebuilt scene graph while we utilize an LLM planner that acts as a high-level controller and participates in the hierarchical scene graph generation process with its commands. While SayPlan [17], MoMa-LLM [20], and SayNav [18] use LLMs and scene graphs in a similar way to our approach, our method distinguishes itself through broader assumptions and a more comprehensive experimental design. SayPlan [17] assumes access to a prebuilt scene graph, whereas we iteratively update the scene graph as the robot explores the environment. MoMa-LLM [20] assumes access to ground-truth perception data, including semantic masks, depth, localization, and handle detection. Similarly, SayNav [18] relies on known robot pose and semantic masks for objects. Although these methods demonstrate real-world robot applications, their reliance on such assumptions limits their deployment in environments where those conditions are unmet. To address this limitation, we develop an end-to-end system that does not rely on these assumptions and integrates our LLM Planner within it.
Some works incorporate different techniques instead of exploiting LLMs with scene graphs for robot navigation tasks. For example, Explore until Confident [96] employ Vision-Language Models (VLMs) to build a semantic map for embodied question answering, replacing frontier-based search by leveraging semantic knowledge of the environment. OK-Robot [97] proposes an end-to-end system for open-ended pick-and-drop tasks using pretrained models. They employ CLIP [37] to convert language queries into semantic vectors, which are then used to search for the closest voxel in a semantic point cloud. PixelNav [95] proposes a direct RGB-based solution that utilizes an LLM planner for waypoint generation in RGB space for object navigation tasks. However, most of these methods either do not create an environment representation that can be used for subsequent tasks in the same environment, or require a prebuilt scene representation.
III Problem Formulation
We address the challenge of developing an autonomous language-driven physical agent capable of operating in dynamic, unknown environments. The agent, a quadruped robot, is equipped with an RGB-D camera, LiDAR, and leg odometry sensors (IMU, joint encoders, and contact sensors). The primary objective is to enable the robot to autonomously search for and navigate to specific objects or rooms in response to natural language queries issued by a remote operator. The robot must operate without prior knowledge of the environment, which consists of multiple interconnected rooms containing various objects. The environment is subject to change over time, with no predetermined types of objects, allowing for arbitrary categories and adding complexity to the search and navigation tasks. The robot must interpret queries, explore the environment, and locate targets in real time while adapting its strategies when initial attempts are unsuccessful. Performance is evaluated based on the robot’s ability to successfully locate and reach queried targets, considering time and path efficiency. The problem encompasses challenges in perception, navigation, natural language understanding, and decision-making in unstructured, real-world settings. The desired outcome is a system that can build and maintain a scalable representation of the environment and objects within it, which can be utilized to handle subsequent queries more efficiently.
IV METHODOLOGY
This section presents our novel framework for addressing the challenges outlined in Section III. Our framework efficiently explores the environment, constructing a hierarchical scene graph that captures object semantics and relationships (Section IV-A1). An LLM-based high-level planner translates human queries into structured commands for execution using the scene graph (Section IV-B). The robot employs safe motion primitives for exploration and navigation, reducing accident risk (Section IV-C). This methodology integrates open-vocabulary semantic mapping, hierarchical scene graph generation, and LLM-based query interpretation with a low-level motion planner and controller, enabling safe autonomous operation in response to the problem outlined in Section III. These modules, as detailed in the following subsections, can be easily substituted as advancements occur in these areas or to accommodate different user needs, provided they adhere to the communication protocols.
IV-A Open-world Hierarchical Scene Graph
IV-A1 Open-vocabulary Semantic Segmentation
The agent uses RGB frames from the RGB-D camera to semantically segment the scene. We use an off-the-shelf open-vocabulary semantic segmentation framework, FC-CLIP [93], that comprises of a class agnostic mask generator, in-vocabulary classifier, and out-of-vocabulary classifier. All modules leverage multi-scale feature maps extracted from a frozen ConvNeXt-L CLIP (Contrastive Language-Image Pretraining) backbone [94]. The mask generator passes these feature maps through a pixel decoder with multi-scale deformable attention, as described in Mask2Former [98], to enhance the pixel features. These enhanced features, along with object query embeddings, are then passed through a series of mask decoders. Each decoder consists of masked cross-attention, self-attention, and a feed-forward network.
Figure 3: Prompt templates used to generate text embeddings for each category. In these templates, every instance of ‘<label>’ is replaced with corresponding category names.
The candidate masks in FC-CLIP require category text embeddings for each class to enable zero-shot classification. These embeddings are generated using a method similar to CLIP, employing a set of predefined template prompts (e.g., ‘a photo of a <label>’, ‘an image of a <label>’) for each class as shown in Fig. 3. For each class, the class name and its synonyms (e.g., cabinet, cupboard, bureau, china cabinet) are individually inserted into these templates, and the resulting texts are processed through a CLIP text encoder. Importantly, each synonym generates its own unique embedding, allowing the model to capture subtle semantic differences. However, during the classification process, these separate embeddings are all associated with a single, general class name (e.g., ‘cabinet’) for output, which aids in LLM room classification and scene graph generation. This approach allows for on-the-fly computation of embeddings for new objects without additional training and enhances the model’s robustness to different object naming conventions. The classification process involves both out-of-vocabulary and in-vocabulary classifiers. The out-of-vocabulary classifier, which applies mask pooling to the frozen CLIP backbone’s feature maps, is designed to maintain CLIP’s open-vocabulary recognition ability for unseen classes. The in-vocabulary classifier, which pools the pixel decoder’s output feature maps, is optimized for better performance on classes seen during training. Both compute classification scores using cosine similarity between the pooled features and each of the synonym embeddings, followed by softmax with a temperature of 0.07. The highest score among synonyms determines the class prediction. These scores from both classifiers are then combined using geometric ensembling, leveraging the strengths of both classifiers to improve overall performance on both seen and unseen classes. Importantly, experiments have shown that using both classifiers also improves performance on out-of-vocabulary classes, demonstrating the synergistic effect of this approach. The mask generator produces 100 fixed mask proposals per image, which are filtered based on classification scores and background class assignment. Remaining masks $\mathcal{M}_{o}$ are associated with their corresponding object class $o$, using the general class name. Objects like ceilings, floors, and walls are segmented but excluded from object mapping to improve robustness against outlier masks. Category text embeddings for COCO Panoptic dataset classes (used to train FC-CLIP) and additional labels relevant to indoor environments, along with their synonyms, are included in the experiments, all mapping to appropriate general class names. This process enables FC-CLIP to perform effective zero-shot segmentation across a wide range of object classes, including those not seen during training, while being flexible to various object naming conventions and providing consistent class names for downstream tasks.
IV-A2 Semantic Object Mapping
The predicted 2D segmentation mask for an object, as described in Section IV-A1 and denoted by $\mathcal{M}_{o}$ (for object $o$) is used to estimate the object’s center $(u_{o},v_{o})$ in the camera frame. These center points are then projected from the camera plane to 3D coordinates $P_{o}^{c}=(X_{o}^{c},Y_{o}^{c},Z_{o}^{c})$ in the camera frame, using depth frame from RGB-D camera, and camera intrinsics. This approach was selected over alternatives such as using 3D bounding boxes or computing the mean of the pointcloud, considering simplicity and computational complexity. These object observations are subsequently converted from the camera frame to the world frame using the pose from the robot’s odometry. We employ an off-the-shelf SLAM toolbox [99] that utilizes the robot’s onboard LiDAR and legged-inertial odometry to build a 2D map of the environment as well as estimate the pose of the robot. The detected objects ($P_{o}^{c}$ in the camera frame) are transformed to the world frame $P_{o}^{w}$ using the estimated pose and associated with previous map objects.
The map objects are ones observed in prior frames and frame objects are objects observed in the current frame. All frame objects observed in the first frame are considered as map objects and if there are multiple map objects belonging to the same class, they are given a different instance number. The instance number is the instance count of a particular class in the scene and is thus initialized at 1 and incremented with new observations. For the subsequent frames, the frame object and map object are associated using density-based clustering (DB-SCAN) [100] based assignment as it offers a good balance between accuracy and computational speed compared to more complex methods like Hungarian matching [101]. In this process, frame objects are projected into the world coordinate frame and DB-SCAN is applied to group frame and map objects of the same class, using a local radius threshold of $\epsilon$ meters. For each resulting cluster, only one map object is retained, while the others are discarded. DBSCAN is effectively able to handle grouping multiple nearby detections, such as when a large object is detected as several smaller bounding boxes across consecutive frames. We consider three cases for grouping objects: for larger objects like tables, doors, and noticeboards, we set $\epsilon=2$ meters to account for their size; for smaller objects, such as books, a reduced threshold of $\epsilon=0.5$ meters is used; and for all other objects, a threshold of $\epsilon=1$ meter is applied. This makes it a good choice to filter out extra bounding boxes corresponding to a single object without using any complex heuristics unlike other filtering methods. All the objects without any associations are assigned as new map objects with unique instance numbers, thus forming a semantic object map $\mathcal{S}$.
The system handles updates in object locations following loop closure and local bundle adjustment using a pose graph-based approach. Each object in the semantic object map $\mathcal{S}$ is associated with the keyframe in which it was first observed. A relative transform $T_{o}^{k}$ between the object and its associated keyframe in the pose graph is maintained. When loop closure or local bundle adjustment occurs and keyframe poses are updated, the location of each object is adjusted by applying the updated keyframe pose to the stored relative transform: $T_{o}^{w}=T_{k}^{w}\cdot T_{o}^{k}$, where $T_{o}^{w}$ is the updated world pose of the object, $T_{k}^{w}$ is the updated pose of the keyframe in the world coordinates, and $T_{o}^{k}$ is the relative transform between the object and its keyframe. An alternative approach involving bundle adjustment using all keyframes that observe an object was considered but not chosen due to its higher complexity and computational cost, favoring simplicity and efficiency in our current system. Our method ensures that object locations remain consistent with the refined map and robot trajectory, maintaining the integrity of the semantic object map even as the underlying SLAM system refines its estimates.
IV-A3 Hierarchical Scene Graph
Indoor environments are, by design, distributed into several semantic components like floors and rooms. Based on the purpose and the objects present in these areas, rooms can be labeled into commonly known language terms. Since these terms are fairly common and consistent across all indoor environments, they become highly informative semantic labels. For e.g., a room labeled as a ‘kitchen’ would have a much higher probability of containing a microwave as opposed to another room labeled as an ‘office’. Hence, as proposed and utilized in previous works [44, 45, 15, 16], a 3D scene graph with room hierarchy is an effective and efficient representation of indoor spaces.
Typically most indoor spaces consist of several rooms connected by walkways called corridors. Hence, it is natural to find multiple clusters of objects while exploring such environments. Making use of this observed property, we propose the classification of the scene into multiple areas, obtained by applying density-based clustering (DB-SCAN [100]) on the coordinates of the map objects. While most previous works rely on language-aligned visual embedding models to find the best match, our experimentation has led us to believe that it is fairly difficult to model the influence of certain combinations of objects toward accurate room classification. Hence, we harness the capabilities of LLMs to generate responses that are consistent with the general notion of such semantic concepts. Providing the labels of all the objects present in each area cluster, we query an LLM to realize the most appropriate room label. To maintain consistency across the environment, we also provide a list of candidate room labels and a few general guidelines in the prompts. To generate these labels in real-time and on-device, we use a memory-efficient 4-bit quantized version of the LLama3 [102] model.
Finally, a tree structure is formed with each node containing properties such as the hierarchy level, text label, location in the map, and parents and children to store the scene graph efficiently (Fig. 2). This type of structure is easily converted to a Python dictionary format that can be directly passed to the LLM-based planner as a prompt.
IV-B LLM Planner
Figure 4: Overview of the system and user prompts for the LLM planner. The system prompt explains the agent’s role, action primitives, and map format. The user prompt provides the map, command history, feedback, and task details. In the initial call to the LLM, command history and feedback are absent, as there is no prior interaction. Feedback includes task status and error messages from previous command executions. ‘<text>’ represents a placeholder for the information summarized in text.
Algorithm 1 Task Execution
1:  Input: user_command
2:  action_history $\leftarrow$ []
3:  feedback $\leftarrow$ None
4:  task_accomplished $\leftarrow$ False
5:  while not task_accomplished do
6:     scene_graph $\leftarrow$ HierarchicalSceneGraph.get_latest()
7:     action $\leftarrow$ LLMPlanner.call(user_command, scene_graph, feedback, action_history)
8:     action_history.append(action)
9:     action_success $\leftarrow$ RobotApi.send(action)
10:     if action_success then
11:        if action = goto(<arg>) then
12:           task_accomplished $\leftarrow$ ask_user(‘Is task done?’)
13:           if not task_accomplished then
14:              feedback $\leftarrow$ ‘Task has not been accomplished.’
15:           else
16:              feedback $\leftarrow$ None
17:           end if
18:        else
19:           task_accomplished $\leftarrow$ False
20:           feedback $\leftarrow$ ‘Task has not been accomplished.’
21:        end if
22:     else
23:        task_accomplished $\leftarrow$ False
24:        feedback $\leftarrow$ RobotApi.get_recent_action_logs()
25:     end if
26:  end while
The LLM Planner uses three action primitives to navigate and interact with the environment: goto(<room name>or <object name>), which navigates the robot to the specified room or near the object and then queries a visual agent to verify if the task is completed; search_room(<room name>), which first navigates the robot to the room given in the function argument and then rotates to detect nearby objects; and explore_globally(), which initiates a frontier-based exploration method. Documentation of these commands, including explanations and the scenarios in which they should be used, is provided to the LLM (see Fig. 4 for prompting details). To ground the LLM in the environment, we supply it with the hierarchical scene graph that stores information about the rooms and the objects within them, generated in parallel by a separate process and provided in JSON format. The LLM selects an action and fills in the arguments using the scene graph to accomplish the user-defined task, while also providing reasoning for its decision. It is also required to provide reasoning for it. The output is directly executed as a function call to the robot API, implemented using the Safe Robot Controller described in Section IV-C.
The plan generation framework operates as follows: If the scene graph is empty, the robot first performs a 360-degree rotation to observe the environment. The LLM is then tasked with generating an action based on the given task and the hierarchical scene graph. After this action is executed, if the task is not yet completed, the updated scene graph is provided, and the LLM is prompted to generate a subsequent action. In case an error occurs—due to incorrect syntax, low-level controller issues, or environmental conditions such as an inaccessible target point—a feedback message explaining the error is generated and provided to the LLM, prompting it to produce another action. Additionally, after the initial action is generated, a command history containing previously executed actions is supplied to prevent the repetition of the same actions. This iterative action generation process is formalized in Algorithm 1. The details of the action primitives provided to the LLM are as follows:
IV-B1 goto(<room name> or <object name>)
This is the navigation primitive used by the LLM to move between rooms and approach objects. It can be used either to go near an object or navigate to a specific room. In the prompts, the Planner is instructed to specify an object name as well as the room it belongs to in the argument when the given task involves navigating to an object. If it is related to navigating to a specific room—like an office with three computers—only the room name is given in the arguments.
This command also allows the LLM Planner to determine whether the task has been completed. After navigating to the object or room specified in the arguments, it prompts a visual agent to verify whether the task has been accomplished. To prevent the Planner from overusing this command after every action, the LLM is conditioned to use it only when it believes the task is complete following the execution of this command. In the proposed framework, user feedback serves as the oracle visual agent. For a fully autonomous system capable of independently determining task completion, a Vision-Language Model (VLM) can be employed.
IV-B2 search_room(<room name>)
This action navigates the robot to the specified room and then performs a 360-degree rotation there. It is designed to be used to explore a partially observed room, allowing the robot to detect objects within and update the room’s graph representation. While this command might be unnecessary in simulated environments with ground truth segmentation, it is crucial in our problem setting. Even if the robot has previously been in the room, imperfect perception may cause it to miss some objects. Therefore, this command provides the Planner with the opportunity to re-observe the room.
IV-B3 explore_globally()
When the scene graph is either empty or lacks sufficient information for the LLM to use the other commands,this primitive is expected to be utilized. It triggers a minute of frontier-based exploration (see Section IV-C3 for details), aiming to map new areas so that the LLM can provide more task-specific commands. After the exploration concludes, the robot performs a 360-degree rotation.
IV-C Safe Robotics Controller
IV-C1 Navigation
The ROS2 navigation stack is used to execute the goto commands from the LLM planner. For the global trajectory planning, we use a wavefront Dijkstra search method, that uses a global cost map computed from the occupancy grid map generated using the LiDAR data. The obstacles in the map are inflated radially, with manually tuned inflation radius and cost scaling factor that determine the motion planning penalty around obstacles. Considering the non-linearity in the considered motion model and the importance of maintaining a planning horizon for improved safety, a sampling-based Model Predictive Path Integral (MPPI) controller is used which outputs the desired velocity commands to the locomotion controller on the robot.
IV-C2 CBF-QP Action Filter
Even though the output of the MPPI controller is aware of the local obstacles through the cost map, it does not guarantee the safety of the actions in terms of colliding with surrounding obstacles. To further improve the safety features of our motion planning pipeline, we adopt the formulation proposed in [103] to introduce a safety filter between the output of the MPPI and the input of the low-level velocity tracker on the robot. This low-level filter operates analogous to the reactive instincts of biological animals, creating a low-latency link between the perception (raw LiDAR pointclouds) and action to filter out dangerous commands that lead to collision.
IV-C3 Exploration
To enable efficient exploration, we use the m-explore ROS2 package [^1] https://github.com/robo-friends/m-explore-ros2 based on [104] with some key modifications. The robot’s current 2D position and the occupancy map generated using SLAM (Section IV-A2) are used. The map has grid points classified as unexplored, occupied, or explored. Unexplored grid points adjacent to explored points are clustered, with the centroid of each cluster defined as a frontier point. Each frontier $\mathcal{F}$ is, therefore, described by its 2D coordinates $(\mathcal{F}_{x},\mathcal{F}_{y})$, distance from the robot $\mathcal{F}_{d}$, number of points $\mathcal{F}_{n}$, and a weight $\mathcal{F}_{w}=\alpha\mathcal{F}_{d}+\mathcal{F}_{n}$. We set $\alpha=50$ to prioritize closer frontiers. The highest-weighted frontier is selected for navigation, and non-navigable frontiers, such as the ones behind obstacles or in narrow spaces, are blacklisted. Unlike the original implementation of m-explore, which halts exploration when all frontiers are either explored or blacklisted, we continue exploration by assigning random goals within already explored areas. This strategy allows the robot to detect new frontiers as it moves closer to frontiers that were previously blacklisted due to measurement errors, making the algorithm more robust.
V EXPERIMENTS
V-A Experiment Setup
The primary goals of our experiments are to demonstrate: (a) the real-world applicability of our pipeline through various end-to-end experiments, (b) the advantages of utilizing a Large Language Model (LLM) as a high-level planner, and (c) the benefits of incorporating hierarchical scene graphs and their integration with the LLM. To achieve these objectives, we structured the experiments into three scenarios: Scenarios A and B focus on object retrieval, where the system locates and navigates to a specific object, and Scenario C addresses room navigation, which involves detecting and navigating to a particular room. We conducted a total of 96 end-to-end experiments, with the detailed setup presented in Table I. In Scenario A (short-range object retrieval), objects were positioned within the same room or area as the robot’s starting point, with no occlusions between the object and the robot. Scenario B (long-range object retrieval) required navigating through corridors and multiple rooms to locate target objects. Additionally, to highlight the benefits of the LLM planner and LLM-based room detection, we performed comparative analyses with baseline methods, as discussed in Section V-C1 and Section V-C2.
Figure 5: Robot Setup: All experiments are conducted using a Unitree Go2 quadrupedal robot equipped with onboard LiDAR, Stereo Camera, and compact computers.
TABLE I: Experiment Scenarios
Task Type Example Prompts No. of Exp.
Scenario A: Short
Range Object Retrieval
find a potted plant, go to a monitor, find a chair 36
Scenario B: Long
Range Object Retrieval
45
Scenario C: Room Navigation find a break room 15
Total 96
All experiments were conducted using a Unitree Go2 quadruped robot as the navigation agent, as depicted in Fig. 5. This system is equipped with a Hesai LiDAR and a Realsense D455 stereo camera mounted on top of the robot, alongside an NVIDIA Jetson Orin NX and an NVIDIA Jetson Orin Nano for efficient and lightweight computation. Additionally, the quadruped robot utilizes its legged-inertial sensors to aid in the estimation of its odometry.
Figure 6: Experiment Environments: The experiments were performed in four distinct environments. Environment A is a large, multi-room space with interconnected corridors. Environment B is a subset of Environment A, while Environment C is a smaller, single-room setting with diverse objects. Environment D comprises two perpendicular corridors, forming an L-shaped layout. The 2D LiDAR maps created using SLAM are shown in the figures.
The experiments were carried out in four office environments featuring multiple rooms and corridors as shown in Fig. 6. These environments contain a variety of typical office objects such as computers, tables, chairs, bags, and potted plants. Furthermore, the environments include various room types such as offices, break rooms, and conference rooms.
Figure 7: Success and failure cases of OrionNav across all experiments. Failure cases are categorized into mapping, exploration, perception, and navigation failures, with detailed breakdowns provided on the right.
Figure 8: Scenario A Examples: The robot observed the queried objects during the first 360-degree rotation and then navigated toward them.
Figure 9: Scenario B Example 1: The system is tasked with locating a bag, which is not visible from the robot’s initial position. The LLM planner issued an exploration command, prompting the robot to explore until the bag is detected, after which the robot successfully navigated to the target.
Figure 10: Scenario B Example 2: The system is tasked with locating a bag. It begins at the starting position (a) and initially searches the office-like space (b). After failing to find the bag, the system initiates global exploration (c, d, e). Upon detecting another office space (f), the system enters and searches it (g), successfully locating the bag.
V-B Results
We present the overall results, including individual failure cases and their corresponding reasons, in Fig. 7. Across a total of 96 trials, we achieved 85 successes, resulting in an overall success rate of over 88%. In two trials, failures occurred due to erroneous SLAM maps, and in one case, a high odometry error caused the system to fail. Excluding these, out of the remaining 93 trials, the exploration algorithm failed to direct the robot to the location containing the queried objects in three cases. From the 90 trials with successful exploration, four trials experienced semantic segmentation errors, where either the object was undetected or misclassified. Additionally, there was one case of failure in the CBF-based navigation system, where the robot collided with an obstacle. Notably, the system did not encounter any failures related to LLM planning or scene graph generation, underscoring their robustness. A detailed breakdown of the success rates for various scenarios is provided in Tab. II, along with the average number of LLM Planner steps. The following subsections offer an in-depth analysis of the system’s performance under different conditions.
TABLE II: Experiment Results: Number of successful trials and the average number of LLM Planner calls
Task Type
No. of
Exp.
No. of
Success
Avg. No. of LLM
Planner steps
Scenario A: Short
Range Object Retrieval
36 32 1.25
Scenario B: Long
Range Object Retrieval
45 41 2.97
Scenario C: Room Navigation 15 12 3.71
Total 96 85
Figure 11: Scenario C Example: The system is tasked to go to an office. The LLM planner issued an exploration command, during which, the robot first encountered a break room. The robot continued exploring until it reached the office, at which point the exploration was halted by the LLM planner, successfully completing the task.
V-B1 Scenario A
In this scenario, both the robot and the target are placed in the same region without any occlusions between them, though the robot is often initially oriented away from the target. We conducted 18 experiments in total. Two objects were selected and placed in three different locations within the environment each, resulting in six distinct target locations. For each target location, we performed three experiments starting from similar positions to account for the stochastic nature of the LLM’s outputs. We observed a high success rate of about 92% in this scenario. During the robot’s initial 360-degree rotation, our framework was able to detect most objects in the environment and assign room labels. If the target object was observed during this rotation, the LLM planner issued the goto command, allowing the robot to navigate directly to the object. In cases where the object was not observed during the first pass, the LLM typically issued a search room command, upon which the object was usually detected.
Three examples of this scenario are demonstrated in Fig. 8, where the system is tasked with finding a bag, a potted plant, and a bottle, respectively. In each case, the system successfully located the queried object during the first 360-degree scan and then navigated toward it. Notably, in all instances, the robot was initially oriented away from the target object.
V-B2 Scenario B
In this scenario, the target object was placed at a larger distance from the robot, often located in separate rooms or corridors. Since the robot lacks prior knowledge of the environment, this scenario demands real-time exploration, making it more challenging than Scenario A. We selected different objects and positioned them at various locations, resulting in a total of 8 distinct target locations. As in Scenario A, we performed three experiments starting from similar positions for each target location to account for the stochastic nature of the LLM’s outputs. Despite the increased complexity, we observed a high success rate of 88%. Unlike in Scenario A, the LLM planner issued global exploration commands at the start. By following these exploration commands, the robot successfully navigated the environment, ultimately detecting the target object. Upon detection, the LLM would issue a goto command, guiding the robot to the target location.
In Fig. 9, we present an example of the robot operating in Scenario B, where the system is tasked with finding a bag. Initially, the bag was not visible from the robot’s starting position. After completing a 360-degree scan, the LLM planner issued a global exploration command, as shown in Fig. 9b. The robot then followed the exploration frontiers and eventually detected the bag upon reaching the position illustrated in Fig. 9c. At this point, the LLM planner issued a goto command, and the robot successfully navigated to the bag. Another example of Scenario B is shown in Fig. 10. This is conducted in the larger ‘Environment B’ from Fig. 6.
V-B3 Scenario C
Scenario C focuses on locating specific rooms, such as an office or a break room. As described in Section IV-A3, our system identifies these rooms using the LLM-based room detection module, which relies on the objects present within the rooms. For this set of experiments, we selected two target rooms and initialized the robot from two different positions for each room. Each experiment was repeated three times, with slight variations in the robot’s starting position. In total, we conducted 12 trials in Scenario C, resulting in only 3 failure cases.
A real-world example of Scenario C is illustrated in Fig. 11, where the system is tasked with finding an office room. Initially, the office room is not visible from the robot’s starting position. As a result, the LLM planner issued an exploration command (Fig. 11b). During exploration, the robot encountered a room containing a coffee maker, microwave, refrigerator, table, and several chairs. The room detection module classified this as a break room, as shown in Fig. 11c. Since the target had not been found yet, the LLM planner instructed the system to continue exploring. The robot then entered a room equipped with multiple computers, monitors, tables, cabinets, and chairs, as depicted in Fig. 11d. Upon identifying this room as an office, the system halted exploration, successfully completing the task.
Figure 12: Comparisons with Baseline Methods: The bar chart illustrates the task completion times (in minutes) for each method across various experiments. Methods that exceed 10 minutes are considered as failures.
V-C Comparative Studies
In this subsection, we compare baseline methods and evaluate the design choices made in the proposed OrionNav, specifically in scenarios involving dynamic environmental changes. The robot is tasked with searching for an object that was initially absent from the environment but added later. For example, this can happen when the object is placed in a room after the robot has already mapped the area. To assess performance, we designed two scenarios: one with a pre-built map (warm start) and another where the robot builds the map from scratch (cold start). In the warm start scenario, the queried object is absent from the pre-built map, as it was placed after mapping was completed. In the cold start scenario, the object was added after the robot had already mapped that specific area. The goal in both cases is to locate the object while minimizing the time required to complete the task. We conducted a total of 9 trials to evaluate our method.
For comparison, we considered two baselines. The first, Frontier-Search, explores the environment by searching through frontiers and transitions to random search once all frontiers are exhausted. This method leverages the ROS2 exploration package, as described in Section IV-C3. To verify object detection in this baseline, the operator manually inspects the object within the object map. The second baseline, Object-Map-Search, utilizes a semantic object map without room labels to locate the target object. For this approach, we introduced a new action primitive, search_object(<object name >), which navigates to the specified object and performs a 360-degree scan to search for other nearby objects. The prompts for the LLM planner were adjusted accordingly.
Figure 13: Example: Dynamically Changes in Environment: The system is provided with a partial map of the environment containing an office room among other locations. A monitor is later placed in the office room and the system is tasked with finding it. In (a), using only the semantic object map without room labels, the system searched near the table in the break room but did not find the monitor. It then explored the environment and eventually located it in the office. In (b), with the proposed hierarchical scene graph, the system identified the office and directly found the monitor inside it.
The comparative results are presented in Fig. 12, where the task completion times are shown as bars for each experiment. Tasks exceeding 10 minutes were deemed failures, indicated by blank bars in the Fig. 12.
V-C1 Frontier-Search
Out of the 9 trials, Frontier-Search failed in 8 attempts and succeeded in only one, completing the task in approximately 3 minutes. In contrast, OrionNav succeeded in all trials with low search times. This superior performance is attributed to the use of the open-world hierarchical scene graph and the LLM-based planner. The LLM Planner identifies the most probable room for the queried object from the scene graph and commands the system to search within that room, leading to successful object detection. In comparison, the Frontier-Search baseline lacks these capabilities, relying solely on frontier-based or random exploration to locate the object. Therefore, even in cases where dynamic changes in the environment prevent the object from being detected during the robot’s initial mapping, OrionNav equips the system with the capability to successfully locate the object in subsequent attempts.
V-C2 Object-Map-Search
The key difference between this baseline and OrionNav is that the baseline uses the semantic object map without room labeling for object search. As a result, its success rate is approximately 56%, compared to OrionNav’s 100% success rate, highlighting the importance of the hierarchical scene graph in the proposed method. An example of this comparison is illustrated in Fig. 13, where the system is given a partial map of an environment that includes an office room, a break room, and a long L-shaped corridor. A monitor is dynamically added to the office room later, and the system is tasked with finding it. As shown in Fig. 13a, the baseline system hypothesizes that a monitor would be located on a table, so it navigates to a nearby table. However, this table was inside the break room and did not have the monitor. The baseline system then issues a global exploration command and eventually reaches the office room, where it detects the monitor. In contrast, OrionNav (Fig. 13b) leverages the hierarchical scene graph to infer that the monitor is likely in an office, navigating directly to it and detecting the monitor in a single step. Object-Map-Search, even in its successful cases, required more time to complete the task compared to OrionNav, as shown in Fig. 12.
VI Conclusion
We proposed OrionNav, an autonomous system that integrates open-vocabulary hierarchical scene graphs with an LLM-based planner for real-time robot autonomy. The system builds a map of the environment using SLAM and fuses open-world semantics, room detection, and LLM-based labeling to generate a hierarchical scene graph. The LLM planner leverages this scene graph, along with user commands, to formulate high-level plans, which are executed by a lower-level safe controller. Extensive experiments in real-world environments demonstrated the system’s high success rates and practical applicability. Future research can explore mobile robots equipped with onboard manipulation capabilities, enabling tasks such as door opening and pick-and-place operations.